OpenStudy (anonymous):

Write this equation in slope-intercept form, and then find the slope and y-intercept. 5y+3x=10

OpenStudy (anonymous):

y=2-(3/5)x slope = -3/5 y-intercept = 2

OpenStudy (anonymous):

How did you get -3? ...

OpenStudy (anonymous):

\[5y+3x=10\]Subtract three x\[5y=10-3x\]\[y=2-\frac{ 3 }{ 5 }x\]

OpenStudy (anonymous):

I don't get what you did after subtracting 3x

OpenStudy (anonymous):

Wait. so you subtracted 3x from 3x and that's out so then you subtracted 3x from 10?

OpenStudy (anonymous):

If you did that, Then you're left with 5y=7x

OpenStudy (anonymous):

I think///

OpenStudy (anonymous):

no, just 5y=10-3x. Ten is not being multiplies by x

OpenStudy (anonymous):

Divide both sides by five

OpenStudy (anonymous):

So I divide 5y by 5 and 3x by 5?

OpenStudy (anonymous):

and ten

OpenStudy (anonymous):

So the answer is y=0.6x+2?

OpenStudy (anonymous):

You dropped the negative sign on three fifths (remember that you subtracted 3x from the right, or added a negative three x)

OpenStudy (anonymous):

Oh//Okay/ Thanks. Another thing,Why am I supposed to add/subtract a -3?

OpenStudy (anonymous):

because slope-intercept form has y by itself on the left: you needed to subtract three x and divide by 5 in order to isolate y.